Mauro Products distributes a single product, a woven basket whose selling price is $12 per unit and whose variable expense is $10 per unit. The company’s monthly fixed expense is $5,400.

Required:

1. Calculate the company’s break-even point in unit sales.

2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.)

3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

Answers

--All working forms part of the answer

A Sale price per unit $12
B Variable cost per unit $10
C = A - B Unit Contribution margin $2
D Fixed Cost $5,400
E = D/C Break even point in unit sales 2700 Answer [1]
F = E x A Break even point in dollar sales $32,400 Answer [2]
A = $ 5400 + 600 New Fixed Cost $6,000
B Unit Contribution margin $2
C = A/B Break even point in unit sales 3000 Answer [3]
D = C x $ 12 Break even point in dollar sales $36,000 Answer [3]